OpenStudy (anonymous):

At the surface (if there is one) of Jupiter the acceleration due to gravity is about 3 times the acceleration due to gravity at the Earth's surface. How much would a 100 kg rock weigh on Jupiter? why?

OpenStudy (anonymous):

it would weigh the same. measurement of mass is irrespective of the gravitational acceleration on it. Its weight changes with the change in gravitational force.

OpenStudy (anonymous):

Wouldn't that be for mass? Not weight?

OpenStudy (anonymous):

idk like gravitational weight or something.

OpenStudy (anonymous):

weight is the product of mass of a body and the gravitaional acceleration on it.

OpenStudy (anonymous):

yes and the acceleration due to gravity is 3x? on jupiter? how would that make it the same weight

OpenStudy (anonymous):

not the same weight. when say, " how much does that weigh?" you mean to talk about the mass

OpenStudy (anonymous):

hmmm i'm unsure. it's for a take home quiz so :/

OpenStudy (anonymous):

@ramkrishna like to say something?

OpenStudy (anonymous):

haha

OpenStudy (anonymous):

@chand Weigh mean not mass it is weigh=mass * acc. due to gravity acc. due to gravity on Jupiter = 9.8*3

OpenStudy (anonymous):

so i would do 9.8*3*100?

OpenStudy (anonymous):

@chand http://dictionary.reference.com/browse/weigh

OpenStudy (anonymous):

@aidanp95 yes.

OpenStudy (anonymous):

are u sur that acceleration gravity on jupiter is greater than on the earth?

OpenStudy (anonymous):

Thank you sirs @ramkrishna @chand

OpenStudy (anonymous):

@aidanp95 owe it to ramakrishna. he corrected me. Sorry for the wrong answer and thanks for correcting me.

OpenStudy (anonymous):

yes. acc. due to gravity on Jupiter is approx. 26

OpenStudy (anonymous):

@chand its okay you tried haha i was confused too

OpenStudy (vincent-lyon.fr):

Let's make things simple: If gE =10 N/kg and gJ = 30 N/kg, then a 100 kg rock would weigh 1000 N on Earth and 3000 N on Jupiter.

OpenStudy (nerdykat):

meow '3'